# Lipschitz-free spaces and Bossard’s reduction argument

### v1

@@ 3. The main results
Let $M$ be a complete separable metric space. By Proposition 3.2, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. By repeating part of the proof above, this time applying Lemma 3.7 to $\mathbb{U}$, we obtain a continuous map ${\Delta}:{\mathbf{Tr}}\to{\mathcal{R}}$ such that $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}$ if $T\in\mathbf{WF}$ and $\mathbb{U}\hookrightarrow M_{\Delta(T)}$ otherwise. By the last part of the proof above, there exists $T\in\mathbf{IF}$ such that $\Delta(T)\in\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$, which implies $\mathbb{U}\hookrightarrow M_{\Delta(T)}\hookrightarrow M$. ∎
Corollary 3.9.
@@
The maps $\Delta$ in the proofs of Theorems 1.1 and 1.2 witness the fact that $\mathbf{WF}$ is reducible to $\mathcal{R}_{\mathrm{dis}}$ and $\mathcal{R}_{\mathrm{p1u}}$. ∎
We finish by making a few remarks about Theorems 1.1 and 1.2. First, it is clear that $M_{\Delta(T)}$ in the proof of Theorem 1.1 is always bounded, so if we require that $X$ is universal for the class of free spaces over countable complete discrete spaces that are in addition bounded, the outcome is the same. Of course, in the proof of Theorem 1.2 we can replace $\mathbb{U}$ by any separable Lipschitz universal space such as $c_{0}$ (by Aharoni’s theorem, see e.g. [benyamini:lindenstrauss:00]*Theorem 7.11). If we add boundedness to the list of properties in Theorem 1.2 then $M$ is Lipschitz universal for all bounded separable metric spaces. This is because we can replace $\mathbb{U}$ in the proof e.g. by any ball in $\mathbb{U}$ (see [vershik:98]*p. 922) or any ball in $c_{0}$.
Finally, we can add the property of being uniformly concave to the lists above and obtain the same conclusions. According to [weaver:18]*Definition 3.33, $(M,d)$ is uniformly concave if, given distinct $x,z\in M$ and $\varepsilon>0$, there exists $\delta>0$ such that $d(x,z)\leqslant d(x,y)+d(y,z)-\delta$ whenever $d(x,y),d(y,z)\geqslant\varepsilon$. Uniformly concave metric spaces are relevant to the extremal structure of free spaces (see e.g. [weaver:18]*Theorem 3.39).
It can be shown that $(\omega,\Delta(T))=M_{\Delta(T)}$ is uniformly concave whenever $T\in\mathbf{WF}$. This holds because $(\omega^{<\omega},\rho)$ is uniformly concave (with $\delta$ independent of $\varepsilon$). To see this we require a quick lemma.
Lemma 3.10.
If $s,t,u\in\omega^{<\omega}$ and $s\wedge u\prec s\wedge t$ then $s\wedge u=t\wedge u$.
@@
As $s\wedge u\preccurlyeq u$ and $s\wedge u\prec s\wedge t\preccurlyeq t$, we have $s\wedge u\preccurlyeq t\wedge u$. Because $s\wedge t,t\wedge u\preccurlyeq t$, they are comparable. However, if $s\wedge t\preccurlyeq t\wedge u\preccurlyeq u$ then $s\wedge t\preccurlyeq s\wedge u$, which is not the case. Thus $t\wedge u\prec s\wedge t\preccurlyeq s$, giving $t\wedge u\preccurlyeq s\wedge u$. ∎
Now let $s,t,u\in\omega^{<\omega}$ be distinct and set $\delta=\min\{2^{-|s|-1},2^{-|u|-1}\}>0$. We will show that $\rho(s,u)\leqslant\rho(s,t)+\rho(t,u)-\delta$, thus demonstrating that $(\omega^{<\omega},\rho)$ is uniformly concave. If $|s\wedge t|\leqslant|s\wedge u|$ then
$\rho(s,t)+\rho(t,u)=2^{-|s\wedge t|-1}+2^{-|t\wedge u|-1}\geqslant 2^{-|s\wedge u|-1}+2^{-|u|-1}\geqslant\rho(s,u)+\delta;$
and we get the same result if $|t\wedge u|\leqslant|s\wedge u|$. On the other hand, if $|s\wedge u|<|s\wedge t|$ then $s\wedge u\prec s\wedge t$ because these elements are comparable, thus $s\wedge u=t\wedge u$ by Lemma 3.10. Then we estimate
$\rho(s,t)+\rho(t,u)=2^{-|s\wedge t|-1}+2^{-|t\wedge u|-1}\geqslant 2^{-|s|-1}+2^{|s\wedge u|-1}\geqslant\delta+\rho(s,u).$
We reach the same conclusion if $|s\wedge u|<|t\wedge u|$.

### v2

@@ 3. Three main results
Let $M$ be a complete separable metric space. By Proposition 3.2, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. Assume that $\mathcal{R}_{\mathrm{dis}}\subseteq\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$. By Lemma 3.10, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}=\mathcal{R}$. By considering e.g. $\mathbb{U}$ or $c_{0}$ (the latter being Lipschitz universal by Aharoni’s theorem, see e.g. [benyamini:lindenstrauss:00]*Theorem 7.11), similarly to above we see that $M$ is also Lipschitz universal. ∎
Proof of Theorem 1.3.
Using [ghawadrah:17], [cuth:doucha:kurka:22a]*Theorem 8 and Proposition 3.1, the set
[…]
is $\mathbf{\Delta}^{1}_{1}$. Let $d^{\prime}\in\mathcal{R}$ and $d\in\mathcal{R}_{\mathrm{BAP}}$ such that $M_{d^{\prime}}$ is isometric to a 1-Lipschitz retract of $M_{d}$. Again by [weaver:18]*Theorems 3.6 and 3.7, $\mathcal{F}({M_{d^{\prime}}})$ is linearly isometric to a 1-complemented subspace of $\mathcal{F}({M_{d}})$ and thus $d^{\prime}\in\mathcal{R}_{\mathrm{BAP}}$. For a contradiction assume $\mathcal{R}_{\mathrm{dis}}\subseteq\mathcal{R}_{\mathrm{BAP}}$. Then $\mathcal{R}_{\mathrm{BAP}}=\mathcal{R}$ by Lemma 3.10. However, this is false because we can pick $d\in\mathcal{R}$ such that $M_{d}$ is isometric to a separable Banach space that fails the AP; as is pointed out in the Introduction, $d\notin\mathcal{R}_{\mathrm{BAP}}$ by [godefroy:kalton:03]*Theorem 3.1.
Now pick $d\in\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$. Again, as stated in the Introduction, $\mathcal{F}({M_{d}})$ has the AP by [aliaga:nous:petitjean:prochazka:21]*Corollary 2.8. That $\mathcal{F}({M_{d}})$ cannot be isomorphic to a dual space is a simple consequence of the classical result of Grothendieck which asserts that a separable Banach space has the MAP whenever it has the AP and is isometric to a dual space. ∎
Remark 3.11.
The set $\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$ is $\mathbf{\Pi}^{1}_{1}$ by Proposition 3.3 and the fact that $\mathcal{R}_{\mathrm{BAP}}$ is $\mathbf{\Delta}^{1}_{1}$. Theorem 1.3 shows that $\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$ is non-empty. By a simple adjustment to the argument we see moreover that $\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$ cannot be $\mathbf{\Delta}^{1}_{1}$: if it was $\mathbf{\Delta}^{1}_{1}$ then $\mathcal{R}_{\mathrm{dis}}\cup\mathcal{R}_{\mathrm{BAP}}$ would be as well, thus it would be equal to $\mathcal{R}$ by Lemma 3.10, which again is false. We make some remarks based on this observation.
  1. (1)
    Given metric spaces $M$ and $N$, let us write $M\stackrel{{\scriptstyle u}}{{\hookrightarrow}}N$ if $M$ is uniformly homeomorphic to a subspace of $N$. Similarly to [clemens:12]*Lemma 4, the set
    $\{(d,d^{\prime})\in\mathcal{R}\times\mathcal{R}\,:\,M_{d}\stackrel{{\scriptstyle u}}{{\hookrightarrow}}M_{d^{\prime}}\}$
    is $\mathbf{\Sigma}^{1}_{1}$ and thus so is $\langle M\rangle^{\mathcal{R},\,u}_{\hookrightarrow}:=\{d\in\mathcal{R}\,:\,M_{d}\stackrel{{\scriptstyle u}}{{\hookrightarrow}}M\}$ whenever $M$ is complete and separable. Evidently $\langle M_{d}\rangle^{\mathcal{R},\,u}_{\hookrightarrow}\subseteq\mathcal{R}_{\mathrm{dis}}$ whenever $d\in\mathcal{R}_{\mathrm{dis}}$. Hence, given a countable subset $C\subseteq\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$, the set $\Gamma:=\bigcup_{d\in C}\langle M_{d}\rangle^{\mathcal{R},\,u}_{\hookrightarrow}\setminus\mathcal{R}_{\mathrm{BAP}}$ is $\mathbf{\Sigma}^{1}_{1}$ and is included in $\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$, which is $\mathbf{\Pi}^{1}_{1}$ but not $\mathbf{\Delta}^{1}_{1}$ from above. By Suslin’s theorem there exists $d^{\prime}\in\mathcal{R}_{\mathrm{dis}}\setminus(\mathcal{R}_{\mathrm{BAP}}\cup\Gamma)$. We can apply transfinite induction to this fact to obtain a sequence $(d_{\alpha})_{\alpha<{\omega_{1}}}$ of elements of $\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$ such that $M_{d_{\beta}}\not\stackrel{{\scriptstyle u}}{{\hookrightarrow}}M_{d_{\alpha}}$ whenever $\alpha<\beta<\omega_{1}$. I don’t know if the $d_{\alpha}$ can be constructed so that $\mathcal{F}({M_{d_{\alpha}}})$ is not isomorphic to $\mathcal{F}({M_{d_{\beta}}})$ whenever $\alpha\neq\beta$.
  2. (2)
    We can replace $\mathcal{R}_{\mathrm{dis}}$ by $\mathcal{R}_{\mathrm{p1u}}$ to show that the set $\mathcal{R}_{\mathrm{p1u}}\setminus\mathcal{R}_{\mathrm{BAP}}$ (which is $\mathbf{\Pi}^{1}_{1}$ by Proposition 3.4) is not $\mathbf{\Delta}^{1}_{1}$. No changes to the proofs above are needed. Given $d\in\mathcal{R}_{\mathrm{p1u}}$, we have $\langle\mathcal{F}({M_{d}})\rangle^{\mathcal{R}}_{\hookrightarrow}\subseteq\mathcal{R}_{\mathrm{p1u}}$ by [aliaga:gartand:petitjean:prochazka:22]*Theorem 4.6. Therefore, by arguing as above, there is a sequence $(d_{\alpha})_{\alpha<\omega_{1}}$ of elements of $\mathcal{R}_{\mathrm{p1u}}\setminus\mathcal{R}_{\mathrm{BAP}}$ such that $\mathcal{F}({M_{d_{\beta}}})\not\hookrightarrow\mathcal{F}({M_{d_{\alpha}}})$ whenever $\alpha<\beta<\omega_{1}$. I don’t know if the spaces $\mathcal{F}({M_{d_{\alpha}}})$ have the AP.
  3. (3)
    While the above results show that there are many examples of complete separable p1u spaces whose free spaces fail the BAP, they don’t provide an explicit description of any of them. It could be instructive to obtain concrete examples.
I am grateful to G. Lancien for suggesting the following problem which Theorem 1.3 leaves open.
Problem 3.12.
Let $M$ be a complete separable p1u metric space. Does there exist a separable Banach space $X$ such that $\mathcal{F}({M})\hookrightarrow X^{*}$?
We finish this section with two direct applications of Lemma 3.9.
Proposition 3.13.
@@
Recall that, by Propositions 3.3 and 3.4, these sets are $\mathbf{\Pi}^{1}_{1}$. Select any complete separable non-p1u metric space $M$. Then the map $\Delta$ furnished by Lemma 3.9 applied to $M$ witnesses the fact that $\mathbf{WF}$ is reducible to $\mathcal{R}_{\mathrm{dis}}$ and $\mathcal{R}_{\mathrm{p1u}}$. ∎
In the final result of this section, we obtain a lower bound on the complexity of the set of metrics in $\mathcal{R}$ whose free spaces satisfy the AP. The complexity calculation is completed in Section 4.
Proposition 3.14.
[…]
@@
Apply Lemma 3.9 to a separable Banach space $M$ that fails the AP. If $T\in\mathbf{WF}$ then $\mathcal{F}({M_{\Delta(T)}})$ has the AP again by [aliaga:nous:petitjean:prochazka:21]*Corollary 2.8. If $T\in\mathbf{IF}$ then $M$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$. Once again by [weaver:18]*Theorems 3.6 and 3.7, $\mathcal{F}({M})$ is linearly isometric to a 1-complemented subspace of $\mathcal{F}({M_{\Delta(T)}})$. Since $\mathcal{F}({M})$ fails the AP, so must $\mathcal{F}({M_{\Delta(T)}})$. Therefore $\Delta$ witnesses the fact that $\mathbf{WF}$ is reducible to $\mathcal{R}_{\mathrm{AP}}$. ∎
4. Complexity and the approximation property
This section is solely concerned with determining the complexity of two classes of Banach spaces having the AP. The proof of the result below is an adaptation of an argument from [ghawadrah:17]. As above, we frame the result in terms of the Polish space $\mathcal{B}$ instead of the Effros-Borel space $\mathcal{SB}$.
Theorem 4.1.
[…]
Proof.
Having in mind Propositions 3.1 and 3.14, it is enough to prove that $\mathcal{B}_{\mathrm{AP}}$ is $\mathbf{\Pi}^{1}_{1}$. Let $(e_{i})_{i\in\omega}\subseteq V$ denote the standard basis of $c_{00}$; then $(e_{i})_{i\in\omega}$ is a linearly independent sequence having dense span in $X_{\mu}$ whenever $\mu\in\mathcal{B}$. Given $R\in\omega$, let $V_{R}=\mathbb{Q}$-$\operatorname{span}(e_{i})_{i<R}\subseteq V$. Define the closed subset
$RC=\left\{{(\mu,x)\in\mathcal{B}\times V^{\omega\times\omega}}\;:\;\,{\mu(x(k,p)-x(k,q))\leqslant 2^{-p}\text{ for all }k,p,q\in\omega,\,q\geqslant p}\right\}$
of the Polish space $\mathcal{B}\times V^{\omega\times\omega}$ (where $V$ has the discrete topology). Then, given finite $F\subseteq\omega$ and $n,k,j\in\omega$, define the closed subset
$TB_{n,F,k,j}=\left\{{(\mu,x)\in\mathcal{B}\times V^{\omega\times\omega}}\;:\;\,{\mu(x(k,q)-x(j,q))\leqslant 2^{-n}\text{ for all }q,n\in\omega,\,q\geqslant n+1}\right\}$
of $\mathcal{B}\times V^{\omega\times\omega}$ and the $\mathbf{\Pi}^{0}_{3}$ subset
$TB=\bigcap_{n\in\omega}\bigcup_{\begin{subarray}{c}F\subseteq\omega\\ \mathrm{finite}\end{subarray}}\bigcap_{k\in\omega}\bigcup_{j\in F}TB_{n,F,k,j}.$
Next, given $n,L,R,N\in\omega$, $v=(v_{i})_{i<N}\in V_{R}^{N}$ and $\alpha=(\alpha_{i})_{i<N}\in\mathbb{Q}^{N}$, define the closed subset
$FR_{L,N,v,\alpha}=\left\{{\mu\in\mathcal{B}}\;:\;\,{\mu\bigg(\sum_{i<N}\alpha_{i}v_{i}\bigg)\leqslant L\mu\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg)}\right\}$
of $\mathcal{B}$, and the closed subset
$AP_{n,L,N,v}=$ $\bigg\{(\mu,x)\in\mathcal{B}\times V^{\omega\times\omega}\,:\,q\leqslant N\text{ and }x(k,q)\in V_{N}\text{ implies}$
$\mu\bigg(x(k,q)-\sum_{i<N}x(k,q)(i)v_{i}\bigg)\leqslant 2^{-n}+(L+2)\cdot 2^{-q}\bigg\}$
of $\mathcal{B}\times V^{\omega\times\omega}$. Then define the $\mathbf{\Pi}^{0}_{5}$ subset
$AP=\bigcap_{n\in\omega}\bigcup_{L,R\in\omega}\bigcap_{N\in\omega}\bigcup_{v\in V_{R}^{N}}\bigg(AP_{n,L,N,v}\cap\bigcap_{\alpha\in\mathbb{Q}^{N}}(FR_{L,N,v,\alpha}\times V^{\omega\times\omega})\bigg).$
of $\mathcal{B}\times V^{\omega\times\omega}$. Now let $\mathcal{A}\subseteq\mathcal{B}$ be the $\mathbf{\Sigma}^{1}_{1}$ image of the projection onto the first coordinate of $(RC\cap TB)\setminus AP\subseteq\mathcal{B}\times V^{\omega\times\omega}$. We show that $\mathcal{B}_{\mathrm{AP}}=\mathcal{B}\setminus\mathcal{A}$ and thus $\mathcal{B}_{\mathrm{AP}}$ is $\mathbf{\Pi}^{1}_{1}$.
Let $\mu\in\mathcal{B}\setminus\mathcal{A}$. We prove that $X_{\mu}$ has the AP, i.e. $\mu\in\mathcal{B}_{\mathrm{AP}}$. Let $K\subseteq X_{\mu}$ be compact and let $\varepsilon>0$. We can assume that $\mu(z)\leqslant 1$ whenever $z\in K$. Let $(z_{k})_{k\in\omega}$ be a dense sequence in $K$. Then pick $x\in V^{\omega\times\omega}$ such that $\mu(x(k,q)-z_{k})\leqslant 2^{-q-1}$ for all $k,q\in\omega$. We verify that $(\mu,x)\in RC\cap TB$. Evidently $(\mu,x)\in RC$. Now let $n\in\omega$. Because $(z_{k})_{k\in\omega}$ is totally bounded, there is finite $F\subseteq\omega$ such that whenever $k\in\omega$, there exists $j\in F$ satisfying $\mu(z_{k}-z_{j})\leqslant 2^{-n-1}$. By the choice of $x$,
$\mu(x(k,q)-x(j,q))\leqslant 2^{-q}+2^{-n-1}\leqslant 2^{-n},\quad q\geqslant n+1.$
It follows that $(\mu,x)\in TB$.
As $\mu\notin\mathcal{A}$ and $(\mu,x)\in RC\cap TB$, it must be that $(\mu,x)\in AP$. Fix $n\in\omega$ such that $2^{-n}\leqslant\varepsilon$. Given $N\in\omega$, by equivalence of norms on finite-dimensional spaces, there exists $\theta_{N}>0$ such that
$\sum_{i<N}|\alpha_{i}|\leqslant\theta_{N}\mu\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg),\quad(\alpha_{i})_{i<N}\in\mathbb{R}^{N}.$ (4.1)
As $(\mu,x)\in AP$, there exist $L,R\in\omega$ and $v^{(N)}\in V_{R}^{N}$, $N\in\omega$, such that
$(\mu,x)\in AP_{n,L,N,v^{(N)}}\cap\bigcap_{\alpha\in\mathbb{Q}^{N}}(FR_{L,N,v^{(N)},\alpha}\times V^{\omega\times\omega}),\quad N\in\omega.$
Given $\alpha\in\mathbb{Q}^{N}$, as $\mu\in FR_{L,N,v^{(N)},\alpha}$,
$\mu\bigg(\sum_{i<N}\alpha_{i}v_{i}^{(N)}\bigg)\leqslant L\mu\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg).$ (4.2)
In particular, $\mu(v^{(N)}_{i})\leqslant L\mu(e_{i})$ whenever $i<N$. Since each $(v_{i}^{(N)})$, $i\in\omega$, is a bounded sequence in $V_{R}\subseteq\operatorname{span}(e_{j})_{j<R}$, by a diagonalisation argument there are $w_{i}\in\operatorname{span}(e_{j})_{j<R}$ and $N_{0}<N_{1}<N_{2}<\ldots$ such that $\mu\big(v^{(N_{\ell})}_{i}-w_{i}\big)\to 0$ for all $i\in\omega$. Using (4.2) and the continuity of $\mu$, given $\alpha\in\mathbb{Q}^{N}$,
$\mu\bigg(\sum_{i<N}\alpha_{i}w_{i}\bigg)\leqslant L\mu\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg).$
Hence we can define a bounded $\mathbb{Q}$-linear operator ${T}:{V}\to{\operatorname{span}(e_{i})_{i<R}}$ by setting $Te_{i}=w_{i}$, $i\in\omega$, and then extending to finite $\mathbb{Q}$-linear combinations of the $e_{i}$. By equicontinuity we extend it again to a bounded linear operator on $X_{\mu}$ satisfying $TX_{\mu}\subseteq\operatorname{span}(e_{i})_{i<R}$ and $\mu(T)\leqslant L$.
By taking a further subsequence of natural numbers, also labelled $(N_{\ell})$, we can ensure that
$\mu\big(v_{i}^{(N_{\ell})}-w_{i}\big)\leqslant\frac{2^{-\ell-1}}{\theta_{\ell}},\quad i<\ell.$ (4.3)
Now let $k,q\in\omega$. Pick $\ell\in\omega$ large enough so that $q\leqslant\ell$ and $x(k,q)\in V_{\ell}$. Then $\ell\leqslant N_{\ell}$ implies $q\leqslant N_{\ell}$ and $x(k,q)\in V_{N_{\ell}}$. Because $(\mu,x)\in AP_{n,L,N_{\ell},v^{(N_{\ell})}}$,
$\mu\bigg(x(k,q)-\sum_{i<N_{\ell}}x(k,q)(i)v_{i}^{(N_{\ell})}\bigg)\leqslant 2^{-n}+(L+2)\cdot 2^{-q}.$ (4.4)
We estimate
$\mu\bigg(\sum_{i<\ell}x(k,q)(i)\big(v_{i}^{(N_{\ell})}-w_{i}\big)\bigg)$ $\leqslant\sum_{i<\ell}|x(k,q)(i)|\mu\big(v_{i}^{(N_{\ell})}-w_{i}\big)$
$\leqslant\frac{2^{-\ell-1}}{\theta_{\ell}}\sum_{i<\ell}|x(k,q)(i)|$ by (4.3)
$\leqslant 2^{-\ell-1}\mu\bigg(\sum_{i<\ell}x(k,q)(i)e_{i}\bigg)$ by (4.1)
$=2^{-\ell-1}\mu(x(k,q))$ as $x(k,q)\in V_{\ell}$
$\leqslant 2^{-\ell}\leqslant 2^{-q}.$
The penultimate inequality above holds because $\mu(x(k,q))\leqslant\mu(z_{k})+2^{-1}\leqslant\frac{3}{2}$ for all $k,q\in\omega$. Therefore
$\mu(x(k,q)-Tx(k,q))$ $=\mu\bigg(x(k,q)-\sum_{i<\ell}x(k,q)(i)w_{i}\bigg)$ as $x(k,q)\in V_{\ell}$
$\leqslant\mu\bigg(x(k,q)-\sum_{i<\ell}x(k,q)(i)v_{i}^{(N_{\ell})}\bigg)+2^{-q}$
$=\mu\bigg(x(k,q)-\sum_{i<N_{\ell}}x(k,q)(i)v_{i}^{(N_{\ell})}\bigg)+2^{-q}$ as $\ell\leqslant N_{\ell}$
$\leqslant 2^{-n}+(L+3)\cdot 2^{-q}$ by (4.4).
Fixing $k\in\omega$ and letting $q\to\infty$ yields $\mu(z_{k}-Tz_{k})\leqslant 2^{-n}$. Since this holds for all $k\in\omega$, by continuity $\mu(z-Tz)\leqslant 2^{-n}\leqslant\varepsilon$ for all $z\in K$. Therefore $X_{\mu}$ has the AP.
Conversely, assume that $\mu\in\mathcal{B}_{\mathrm{AP}}$. Let $(\mu,x)\in RC\cap TB$. We show that $(\mu,x)\in AP$ and hence $\mu\in\mathcal{B}\setminus\mathcal{A}$. Let $n\in\omega$. Given $k\in\omega$, let $z_{k}=\lim_{q}x(k,q)\in X_{\mu}$. As $(\mu,x)\in TB$ as well, it is easy to see that $\left\{{z_{k}}\;:\;\,{k\in\omega}\right\}$ is totally bounded and thus
$K:=\overline{{\left\{{z_{k}}\;:\;\,{k\in\omega}\right\}}}\subseteq X_{\mu}$
is compact. As $X_{\mu}$ has the AP, there exists a bounded linear finite-rank operator ${T}:{X_{\mu}}\to{X_{\mu}}$ such that $\mu(z-Tz)<2^{-n}$ for all $z\in K$. Fix $L\in\omega$ such that $\mu(T)<L$. By a perturbation argument, we can assume there exists $R\in\omega$ such that $TX_{\mu}\subseteq\operatorname{span}(e_{i})_{i<R}$. Let $B>0$ such that $\mu(z)\leqslant B$ for all $z\in K$. Given $N\in\omega$, pick $\theta_{N}>0$ to satisfy (4.1) above and choose $v\in V^{N}_{R}$ such that
$\mu(v_{i}-Te_{i})\leqslant\frac{1}{\theta_{N}}\min\bigg\{\frac{2^{-N}}{B+\frac{1}{2}},L-\mu(T)\bigg\},\quad i<N.$ (4.5)
Then, given $\alpha\in\mathbb{Q}^{N}$,
$\mu\bigg(\sum_{i<N}\alpha_{i}v_{i}\bigg)$ $\leqslant\mu\bigg(\sum_{i<N}\alpha_{i}Te_{i}\bigg)+\sum_{i<N}|\alpha_{i}|\mu(v_{i}-Te_{i})$
$\leqslant\mu(T)\mu\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg)+(L-\mu(T))\mu\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg)$ by (4.1), (4.5)
$=L\bigg(\sum_{i<N}\alpha_{i}e_{i}\bigg).$
We conclude that
$(\mu,x)\in\bigcap_{\alpha\in\mathbb{Q}^{N}}(FR_{L,N,v,\alpha}\times V^{\omega\times\omega}).$
We also need to show that $(\mu,x)\in AP_{n,L,N,v}$. Assume that $q\leqslant N$ and $x(k,q)\in V_{N}$. As $(\mu,x)\in RC$ and $\mu(T)\leqslant L$, we have
$\mu(x(k,q)-z_{k})\leqslant 2^{-q}\quad\text{and}\quad\mu(Tx(k,q)-Tz_{k})\leqslant L\cdot 2^{-q}.$ (4.6)
Moreover,
$\mu\bigg(Tx(k,q)-\sum_{i<N}x(k,q)(i)v_{i}\bigg)$ $\leqslant\sum_{i<N}|x(k,q)(i)|\mu(Te_{i}-v_{i})$ as $x(k,q)\in V_{N}$
$\leqslant\frac{2^{-N}}{B+\frac{1}{2}}\mu\bigg(\sum_{i<N}x(k,q)(i)e_{i}\bigg)$ by (4.1), (4.5)
$=\frac{2^{-N}}{B+\frac{1}{2}}\mu(x(k,q))$
$\leqslant 2^{-N}\leqslant 2^{-q}.$
The penultimate inequality above holds as $\mu(x(k,q))\leqslant\mu(z_{k})+2^{-1}\leqslant B+\frac{1}{2}$. Given that $\mu(z_{k}-Tz_{k})\leqslant 2^{-n}$, by (4.6) and the above we conclude that
$\mu\bigg(x(k,q)-\sum_{i<N}x(k,q)(i)v_{i}\bigg)\leqslant 2^{-n}+(L+2)\cdot 2^{-q}.$
Therefore $(\mu,x)\in AP_{n,L,N,v}$. The proof is complete. ∎